OpenStudy (anonymous):

Solve for r. 5r − 3(2r + 5) = 9

OpenStudy (anonymous):

7

OpenStudy (anonymous):

SHOW WORK

OpenStudy (anonymous):

SHOW YOUR WORK BAC JONES

OpenStudy (anonymous):

5r − 3(2r + 5) = 9 5r - 6r + 15 = 9 -1r + 15 = 9 15 - 9 = r r = 6

OpenStudy (anonymous):

\[5r − 3(2r + 5) = 9\] \[5r-6r-15=9\] \[-r-15=9\] \[-r=-6\] \[r=6\]